the demand for tylenol, a chemical sold in a perfectly competitive market, is given by Q = 200 - 2P, where the quantity is measured in pounds. the private marginal cost is MC = 20 + 0.75Q, but producing this chemical also leads to an external marginal damage of MD = 5 + 0.25Q.

What is the equilibrium price and quantity of this good if the externality is ignored?
Answer #1

Q= 200-2P

2P= 200-Q

P= 100 - 0.5Q

When externality is ignored, P= MC

20 + 0.75Q= 100 - 0.5Q

1.25Q= 80

Q= 64

P= 100-0.5(64)= 68